OpenStudy (anonymous):

Please help with one question! will Fan and medal! Skip designs tracks for amusement park rides. For a new design, the track will be elliptical. If the ellipse is places on a large coordinate grid with its center at (0,0), the equation x^2/2500+y^2/8100=1 models the path of the track. The units are given in yards. How long is the major axis of the track? Explain how you found the distance.

OpenStudy (anonymous):

90

OpenStudy (anonymous):

Can you please explain how you got that answer?

OpenStudy (anonymous):

@eHoaX

OpenStudy (anonymous):

anyone?

OpenStudy (anonymous):

look at these 2 pics from Mr.google

OpenStudy (anonymous):

in the first one h and k are both zero's because the center of the elipse is (0,0) === the major axis is the length of the widest part of the elipse, here is is on the x-axis, sometimes depending how they draw the elipse it could be on the why axis === in your eliptical formula what is the value of "a" and "b" if we wrote like this - given that h and k are zeros because the center of the elipse is 0,0 ? \[\frac{ x ^{2} }{ a ^{2} } + \frac{ y ^{2} }{ b ^{2} } = 1\]

OpenStudy (anonymous):

@ganeshie8 hey buddy, see u r online :) , can I please ask u to verify the above answer if u got some time...please?

OpenStudy (anonymous):

thanks @ganeshie8 ;)

ganeshie8 (ganeshie8):

lol this notation stuff of a and b is very tricky.. major axis is always the longest segment : a = 90 since 90 > 50 so major axis = 2a = 2*90

ganeshie8 (ganeshie8):

its a vertical ellipse wid major axis along y-axis, so the equation wud be : \(\large \dfrac{x^2}{b^2} + \dfrac{y^2}{a^2} = 1\) this is valid if we fix our notations for \(\large a\) to be major axis..

ganeshie8 (ganeshie8):

i always get confuse between a and b, i think the simple way to remember is : long = major short = minor

OpenStudy (anonymous):

DAMN!!! u r 100% correct @ganeshie8 I got caught out here, thanks buddy I'll delete my above wrong answer so the answer is 180 , double what @eHoaX had, he was also on the right track ...almost..